# All-Sky Kinematics of the Distant Halo: The Reflex Response to the LMC

### v1

@@ Appendix A All-Sky Velocity Maps vs Distance
Appendix B Posterior Distribution of Distance-Dependent Reflex Model Parameters
Figure 14 shows the posterior distributions of all 18 parameters from our distance-dependent $v_{\mathrm{travel}}$ continuity model, described in §$\S$4. These posterior distributions are quantitatively summarized in Table 1
[…]
Appendix C Fitting the PP21 Dataset
[…]
Furthermore, we directly fit the northern hemisphere K giant dataset [^3] https://github.com/michael-petersen/ReflexMotion from PP21 using our distance-dependent continuity model from §$\S$4. The direction of the reflex motion inferred from our fit to the PP21 dataset is consistent with the results in PP21 across all distances. Furthermore, the amplitude of the measured reflex motion is $\approx 30$ km s-1, likewise consistent with PP21. We conclude that our modeling framework is consistent with that of PP21 and YPP24. The difference between our measurement and theirs is therefore likely due to differences in the underlying datasets — chiefly our inclusion of the MagE stars observed from the southern hemisphere, and more distant H3 data in the northern hemisphere.
Appendix D Distance-Independent Reflex Model

### v2

@@ Appendix D All-Sky Velocity Maps vs Distance
Appendix E Distance-Independent Reflex Model
@@
Table 4: Summary of the best-fit parameters from our $v_{\mathrm{travel}}$ model that uses all stars from $40-120$ kpc.
         Parameter          Value
$v_{\mathrm{travel}}$ [km s-1] $17^{+4}_{-4}$
$\ell_{\mathrm{apex}}$ [deg] $-86^{+35}_{-49}$
$b_{\mathrm{apex}}$ [deg] $-73^{+7}_{-10}$
$\langle v_{\mathrm{r}}\rangle$ [km s-1] $-16^{+3}_{-3}$
$\langle v_{\mathrm{\phi}}\rangle$ [km s-1] $6^{+2}_{-2}$
$\langle v_{\mathrm{\theta}}\rangle$ [km s-1] $1^{+4}_{-4}$
${\sigma_{\mathrm{v,r}}}$ [km s-1] $95^{+2}_{-2}$
${\sigma_{\mathrm{v,\ell}}}$ [km s-1] $58^{+2}_{-2}$
${\sigma_{\mathrm{v,b}}}$ [km s-1] $50^{+1}_{-1}$
@@
Appendix F Posterior Distribution of Distance-Dependent Reflex Model Parameters
Figure 21 shows the posterior distributions of all 18 parameters from our distance-dependent $v_{\mathrm{travel}}$ continuity model, described in §$\S$4. These posterior distributions are quantitatively summarized in Table 1.
Figure 22 shows the information that is missing from Figure 21, namely the covariance of each parameter at $r_{\mathrm{gal}}=40$ kpc versus its counterpart at $r_{\mathrm{gal}}=120$ kpc. There is modest correlation in all panels, which is to be expected for the same physical parameter at different distances. However, all parameters are well-constrained relative to their prior distributions at both distances, confirming that the linear continuity model is well-measured over the distance range considered.
[…]
Figure 22: Posterior distributions for all parameters in the $v_{\mathrm{travel}}$ continuity model, comparing the covariance between each parameter at $r_{\mathrm{gal}}=40$ kpc and $r_{\mathrm{gal}}=120$ kpc. The axes are scaled to span the prior range of each parameter, and the axis limits are indicated in each panel.